### v5

@@ Appendix A Auxiliary concept definitions
Appendix B Theoretical relations with the persistent homology theory and related concepts
@@
In Eq. (39), matrix $\mathbf{Z}^{\left(q\right)}$ is defined as $\mathbf{Z}^{\left(q\right)}=\left[\mathbf{F}^{\left(q+1\right)}\right]^{\top}$, where boundary matrix, $\mathbf{F}^{\left(q+1\right)}$, and its transpose, $\left[\mathbf{F}^{\left(q+1\right)}\right]^{\top}$, are the discrete counterparts of $\partial^{\left(q+1\right)}$ and $\partial^{\left(q+1,*\right)}$ such that the combinatorial Laplacian matrix in Eqs. (37-38) can be expressed as $\mathbf{L}^{\left(q\right)}_{C}=\mathbf{F}^{\left(q+1\right)}\left[\mathbf{F}^{\left(q+1\right)}\right]^{\top}+\left[\mathbf{F}^{\left(q\right)}\right]^{\top}\mathbf{F}^{\left(q\right)}$. Matrix $\mathbf{S}^{\left(q\right)}$ is a diagonal matrix that discretizes the Hodge star [60] on simplicial complexes. The Hodge star is an operator that transforms $q$-forms to $\left(h-q\right)$-forms in differential geometry theories [60]. Meanwhile, the original Hodge Laplacian, $\Delta^{\left(q\right)}_{D}$, can be approximated by a discrete counterpart, $\left[\mathbf{S}^{\left(q\right)}\right]^{-1}\mathbf{L}^{\left(q\right)}_{C}$ [60].
@@
Our idea arises from a simple fact. No matter we implement a SRG framework with the multi-order Laplacian or the high-order path Laplacian, when this SRG renormalizes the system on a $p$-order according to the properties of $q$-order interactions ($p\leq q$), we can always record the set of units in $\mathbf{G}^{\left(p\right)}_{k}$ that are aggregated into each super-unit in $\mathbf{G}^{\left(p\right)}_{k+1}$. By repeating this recording in every iteration, we can recursively find the set of initial units in $\mathbf{G}^{\left(p\right)}_{1}$ aggregated into each super-unit in $\mathbf{G}^{\left(p\right)}_{k+1}$ for each $k\geq 1$. In other words, we can know how the initial structure formed by system units is changed by the SRG after $k$ iterations. If we set $p=1$ and search simplices in the initial system, $\mathbf{G}^{\left(1\right)}_{1}$, we can use the recursively tracked information to identify how many simplices in $\mathbf{G}^{\left(1\right)}_{1}$ are added, removed, or modified by the SRG after $k$ iterations. Therefore, similar to the filtration process in the persistent homology theory [62, 76], the SRG also offers an opportunity to study the lifetime of different topological properties (e.g., the Betti number). Given a simplicial complex $K$, the SRG also creates a filtration sequence of sub-complexes $\left(K_{k}\right)_{k=0}^{m}$ (here $m$ denotes the number of iterations in the SRG). Certainly, the SRG is different from the common filtration processes that persistently add or remove simplices following deterministic rules [62, 76]. It has more intricate effects on the topological properties of the system.
Appendix C The SRG in a form of conventional renormalization group theories
@@
where $\left[\mathbf{L}^{\left(q\right)}\right]_{ij}$ can be selected as $\left[\mathbf{L}^{\left(q\right)}_{M}\right]_{ij}$ or $\left[\mathbf{L}^{\left(q\right)}_{H}\right]_{ij}$. We refer to $\mathcal{H}^{\left(q\right)}$ as the $q$-order functional Hamiltonian, which characterizes the system behaviours on the $q$-order. If we write $\mathbf{K}_{ij}^{\left(q\right)}=\left[\mathbf{L}^{\left(q\right)}\right]_{ij}+r\delta_{ij}$, we notice that $\mathbf{K}_{ij}^{\left(q\right)}$ is essentially equivalent to the term of $\phi^{4}$ theory analyzed in Ref. [31]. The only difference lies in that the PCA-like RG proposed by Ref. [31] does not study high-order interactions. Given this connection, we can have following ideas after learning Ref. [31]:
@@
      As shown by the perturbation analysis of the PCA-like RG [31], analyzing a model with Gaussian fluctuations does not un-conditionally imply a trivial consequence of the central limit theorem. Although a RG is initially proposed for a translation invariant system with local interactions, it may also be applicable to more complex systems and identify non-Gaussian behaviours [31]. For instance, the initial state to renormalize in diverse complex systems can be non-Gaussian. Meanwhile, the departures from a Gaussian state can become either more or less important under the scale transformation by the RG. These conditions are all possible to create non-trivial phenomena.
@@
      Most importantly, perhaps a frequently neglected point is that $\mathbf{K}^{\left(q\right)}$ can be defined with $q>1$ (in fact, this is the most common case in the SRG). Under this condition, the remaining terms in Eq. (44) can not be treated as the descriptions of conventional pairwise couplings. Instead, they have implicitly include $q$-order interactions among units, which require the simultaneous participation of more than two units when $q>1$. For instance, although $\mathbf{K}_{ij}^{\left(4\right)}$ focuses on the $4$-order relation between units $i$ and $j$, the relation itself actually corresponds to the $4$-order interactions involved with unit $i$, unit $j$, and three other units, which can happen over a large region and are beyond pairwise local interactions. Consequently, by considering a high-order functional Hamiltonian, we can study more complex correlations among units, providing opportunities for non-trivial fixed points to occur in the SRG.
Appendix D The behaviour of the specific heat

### v7

@@ Appendix A Auxiliary concept definitions
Appendix B Theoretical backgrounds of high-order Laplacian operators
Given a system with diverse orders of interactions, we can treat it as a simplicial complex, $K$, following Appendix A. In this simplicial complex, we can search a $n$-simplex to verify its potential existence (here $n+1$ is no more than the size of the system). By assigning the identified simplices with different physics meanings, we can use them to model various types of high-order interactions. In our work, we primarily focus on two kinds of basic high-order interactions. The first type of $n$-order interactions occur if and only if the associated $n+1$ units are simultaneously engaged in (i.e., the interaction becomes ineffective unless all the $n+1$ units participate in it at the same time). The second type of $n$-order interactions refer to the situations where $n+1$ units form an irreducible sequence of participation (i.e., the interaction becomes ineffective unless these $n+1$ units participate in it following a sequence). In this section, we elaborate the high-order Laplacian operators associated with them.
Multi-order Laplacian operator
We begin with the first type of $n$-order interactions that manifest as the simultaneous participation of $n+1$ related units. We expect to see a $n$-simplex, $\sigma^{\left(n\right)}$, if there exists a certain $n$-order interaction. This $n$-simplex is observable if we analyze the system on $h$-order ($h\geq n$). The corresponding Laplacian operator is referred to as the multi-order Laplacian operator, $\mathbf{L}_{M}^{\left(n\right)}$, which is originally proposed by Ref. [9]. To define such an operator, we actually need to measure the difference between the following numbers
$\left[\mathbf{L}_{M}^{\left(n\right)}\right]_{ij}=\delta_{ij}\cdot\text{the number of }\sigma^{\left(n\right)}\text{ containing unit }v_{i}-\text{the number of }\sigma^{\left(n\right)}\text{ containing units }v_{i}\text{ and }v_{j},$ (34)
where $\delta_{ij}$ denotes the Kronecker delta function. The non-triviality is that there may exist diverse simplices in the system and $\sigma^{\left(n\right)}$ can be a face of multiple simplices. Therefore, a unit can participate in many $n$-order interactions. Two coupled units do not necessarily share only one $n$-order interaction. In the most general situation, we need to go through all the possibilities to count the numbers mentioned in Eq. (34).
To search these possibilities, we let $\{i\}_{n+1}\subset V$ be an arbitrary set of $n+1$ units that contains unit $v_{i}$, where $V$ is the set of all units. Similarly, we denote $\{i,j\}_{n+1}\subset V$ as the set of $n+1$ units where units $v_{i}$ and $v_{j}$ are included. Given a set $\{i\}_{n+1}$ during the search process, we check whether the units contained by it form a $n$-simplex or not. For clarity, we denote $I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right)=1$ if there is a $n$-simplex, $\sigma^{\left(n\right)}$, formed by these units and define $I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right)=0$ if there is no such $n$-simplex. Based on these definitions, we can count the number of $n$-simplices containing units $v_{i}$ and $v_{j}$ as
$\mathbf{A}_{ij}^{\left(n\right)}=(1-\delta_{ij})\sum_{\{i,j\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right).$ (35)
A non-trivial property of Eq. (35) is that summing Eq. (35) over all possible units $v_{j}$ does not directly obtain the number of $n$-simplices containing unit $v_{i}$ unless $n=1$ (i.e., for pairwise interactions)
$\sum_{j}\mathbf{A}_{ij}^{\left(n\right)}$ $=\sum_{j}\left(1-\delta_{ij}\right)\sum_{\{i,j\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right),$ (36)
$=\sum_{j\neq i}\sum_{\{i,j\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right),$ (37)
$=n\sum_{\{i\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right).$ (38)
We obtain Eq. (38) from Eq. (37) because we repeatedly count each $n$-simplex $n$ times while summing $I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right)$ over all units $v_{j}$ in Eq. (37). As a simple instance, let us consider a case with $n=3$. We assume that $v_{i}$, $v_{j}$, $v_{k}$, and $v_{l}$ form a $3$-simplex, $\sigma^{\left(3\right)}$. This simplex is considered when we analyze $I_{\sigma^{\left(3\right)}}\left(\{i,j\}_{4}\right)$, $I_{\sigma^{\left(3\right)}}\left(\{i,k\}_{4}\right)$, and $I_{\sigma^{\left(3\right)}}\left(\{i,l\}_{4}\right)$. Thus, it is repeatedly counted $3$ times. There is no repetition for pairwise interactions (i.e., $n=1$) because a pairwise interaction only consists of unit $v_{i}$ and another one unit, which is analyzed once.
In Eq. (38), the term scaled by factor $n$ equals the number of $n$-simplices containing unit $v_{i}$. Therefore, we have
$\sum_{j}\mathbf{A}_{ij}^{\left(n\right)}$ $=n\mathbf{D}_{i}^{\left(n\right)},$ (39)
where $\mathbf{D}_{i}^{\left(n\right)}$ counts the $n$-simplices that include unit $v_{i}$. Eq. (39) directly leads to
A remaining question is how to determine the values of $I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right)$ and $I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right)$ in Eqs. (35-38). For convenience, we define $S^{\{i\}}_{n+1}$ as the set of all permutations on $\{i\}_{n+1}$. Because the $n$-simplex studied in our work is a clique of $n+1$ units, we can calculate the indicator function as
$I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right)=\prod_{\left(k_{1}\ldots k_{n}\right)\in S^{\{i\}}_{n+1}}\mathbf{A}_{k_{1}k_{2}}\cdots\mathbf{A}_{k_{n-1}k_{n}},$ (41)
where $\left(k_{1}\ldots k_{n}\right)$ denotes a permutation $\mathbf{A}=\mathbf{A}^{\left(1\right)}$ is the classic adjacency matrix of units (i.e., this matrix can be derived on pairwise interactions). According to the recursive attribute of a simplex mentioned in Appendix A (i.e., each face of a simplex is also a simplex), the units in $\{i\}_{n+1}$ can form a $n$-simplex if and only if each sub-set of $\{i\}_{n+1}$ that contains $m+1$ units also forms a $m$-simplex ($m\leq n$). By going through all permutations in $S^{\{i\}}_{n+1}$, we can check the existence of all possible $m$-simplices. Therefore, starting from pairwise interactions, Eq. (41) enables us to verify the existence of $2$-order, $3$-order, and eventually $n$-order interactions. For simplicity, we mark $\omega=\left(k_{1}\ldots k_{n}\right)$ and define
$\mathbf{M}_{\omega}=\mathbf{A}_{k_{1}k_{2}}\cdots\mathbf{A}_{k_{n-1}k_{n}},$ (42)
which supports us to write Eq. (42) as $I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right)=\prod_{\omega\in S^{\{i\}}_{n+1}}\mathbf{M}_{\omega}$.
High-order path Laplacian operator
Then, we formulate the second type of $n$-order interactions, which manifest as the sequential actions of $n+1$ related units after they participate in the interplay progressively. The second type of $n$-order interactions differ from the first type because the participation of units happen sequentially rather than simultaneously. In both types of high-order interactions, none of the involved units is dispensable. Therefore, we need to analyze both the faces of simplices (i.e., to ensure that all related units are involved) and the paths defined on these faces (i.e., to describe the sequential participation) to model the second type of $n$-order interactions.
The Laplacian operator corresponding to the second type of $n$-order interactions is referred to as the high-order path Laplacian operator, whose general form is defined as
$\left[\mathbf{L}_{H}^{\left(n\right)}\right]_{ij}=$ $\frac{1}{n}\big{(}\delta_{ij}\cdot\text{the number of paths that traverse all units on }\sigma^{\left(n\right)}\text{ without repetition and terminate at }v_{i}$
[…]
Here a path can be also treated as a sequence of $1$-simplices. The factor $\frac{1}{n}$ is used to normalize the propagation speed of interactions along paths (i.e., there are $n$ end-to-end $1$-simplices in each path).
In Eq. (43), the number of paths that traverse all units on a $n$-simplex without repetition, initiate at unit $v_{j}$, and terminate at unit $v_{i}$ can be calculated as
$\mathbf{B}_{ij}^{\left(n\right)}=\left(1-\delta_{ij}\right)\sum_{\{i,j\}_{n+1}\subset V}\left(n-1\right)!I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right),$ (44)
where the coefficient $\left(n-1\right)!$ arises from the fact that the rest part of units in $\{i,j\}_{n+1}$ can form $\left(n-1\right)!$ kinds of unique sequences after we fix units $v_{i}$ and $v_{j}$ as the endpoints of these sequences. Based on simple derivations, we have
$\mathbf{B}_{ij}^{\left(n\right)}=\left(n-1\right)!\mathbf{A}_{ij}^{\left(n\right)},$ (45)
which indicates the difference in freedom degrees between the first and the second type of $n$-order interactions. Specifically, if units $v_{i}$ and $v_{j}$ share a single $n$-order interaction of the first type, there are $\left(n-1\right)!$ possibilities for the second type of $n$-order interactions to occur and include units $v_{i}$ and $v_{j}$.
Another difference between the first and the second type of $n$-order interactions is that summing Eq. (44) over all possible units $v_{j}$ directly leads to the number of paths that traverse all units on a $n$-simplex without repetition and terminate at unit $v_{i}$, irrespective of whether these interactions are pairwise ($n=1$) or high-order ($n>1$)
$\sum_{j}\mathbf{B}_{ij}^{\left(n\right)}$ $=\sum_{j}\left(1-\delta_{ij}\right)\sum_{\{i,j\}_{n+1}\subset V}\left(n-1\right)!I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right),$ (46)
$=\left(n-1\right)!\sum_{j\neq i}\sum_{\{i,j\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i,j\}_{n+1}\right),$ (47)
$=\left(n-1\right)!\left[n\sum_{\{i\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right)\right],$ (48)
$=n!\sum_{\{i\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right).$ (49)
Eq. (48) is derived from Eq. (47) using Eqs. (37-38). We can discover that Eq. (49) equals to the number of paths that traverse all units on a $n$-simplex without repetition and terminate at unit $v_{i}$. This is because the rest part of units in $\{i\}_{n+1}$ can form $n!$ unique sequences if we fix unit $v_{i}$ as the endpoint. For convenience, we define $\mathbf{P}_{i}^{\left(n\right)}=n!\sum_{\{i\}_{n+1}\subset V}I_{\sigma^{\left(n\right)}}\left(\{i\}_{n+1}\right)$ as the number these paths and write Eq. (50) as
According to Eqs. (38-39) and Eq. (49), we can see that
$\mathbf{P}_{i}^{\left(n\right)}=n!\mathbf{D}_{i}^{\left(n\right)}.$ (51)
Differences between two kinds of high-order interactions
Based on Eq. (45) and Eq. (51), the multi-order Laplacian operator and the high-order path Laplacian can be related by
$\left[\mathbf{L}_{H}^{\left(n\right)}\right]_{ij}$ $=\frac{1}{n}\left(\delta_{ij}n!\mathbf{D}_{i}^{\left(n\right)}-\left(n-1\right)!\mathbf{A}_{ij}^{\left(n\right)}\right),$ (52)
$=\frac{\left(n-1\right)!}{n}\left[\mathbf{L}_{M}^{\left(n\right)}\right]_{ij}.$ (53)
Please note that the factorial satisfies $0!=1$. Based on Eq. (53), we can see that the joint effects of the first type of $n$-order interactions (i.e., simultaneous participation) are no less than the second type of $n$-order interactions (i.e., sequential participation) only when $n\leq 2$. Specifically, we can consider the evolution of system dynamics mentioned in Sec. II. We assume that systems units follow the sequential participation
$\mathbf{x}=$ $\exp\left(-\tau\mathbf{L}_{H}^{\left(n\right)}\right)\mathbf{x}_{0}.$ (54)
We calculate $\mathbf{L}_{M}^{\left(n\right)}=\sum_{\lambda_{M}^{\left(n\right)}}\lambda_{M}^{\left(n\right)}\big{|}\lambda_{M}^{\left(n\right)}\big{\rangle}\big{\langle}\lambda_{M}^{\left(n\right)}\big{|}$, where each $\lambda_{M}^{\left(n\right)}$ is an eigenvalue of $\mathbf{L}_{M}^{\left(n\right)}$. Then, we can reformulate Eq. (54) as
$\mathbf{x}$ $=\exp\left(-\tau\frac{\left(n-1\right)!}{n}\mathbf{L}_{M}^{\left(n\right)}\right)\mathbf{x}_{0},$ (55)
$=\left(\sum_{\lambda^{\left(n\right)}}\exp\left(-\tau\frac{\left(n-1\right)!}{n}\lambda_{M}^{\left(n\right)}\right)\big{|}\lambda_{M}^{\left(n\right)}\big{\rangle}\big{\langle}\lambda_{M}^{\left(n\right)}\big{|}\right)\mathbf{x}_{0}$ (56)
using the property of matrix exponential. Based on Eq. (56), we can see that the decay rate of dynamic evolution, which is measured by $\tau\frac{\left(n-1\right)!}{n}\lambda_{M}^{\left(n\right)}$, depends on the type of $n$-order interactions and the actual value of $n$. When $n=1$ (i.e., pairwise interactions), the decay rate maintains the same no matter units follow the simultaneous or the sequential participation since $\frac{\left(n-1\right)!}{n}=1$. When $n=2$, the sequential participation leads to a smaller decay rate compared with the simultaneous participation since $\frac{\left(n-1\right)!}{n}<1$. When $n>2$, the sequential participation causes a larger decay rate since $\frac{\left(n-1\right)!}{n}>1$.
@@ Appendix B Theoretical backgrounds of high-order Laplacian operators
Appendix C Theoretical relations with the persistent homology theory and related concepts
@@
In Eq. (64), matrix $\mathbf{Z}^{\left(q\right)}$ is defined as $\mathbf{Z}^{\left(q\right)}=\left[\mathbf{F}^{\left(q+1\right)}\right]^{\top}$, where boundary matrix, $\mathbf{F}^{\left(q+1\right)}$, and its transpose, $\left[\mathbf{F}^{\left(q+1\right)}\right]^{\top}$, are the discrete counterparts of $\partial^{\left(q+1\right)}$ and $\partial^{\left(q+1,*\right)}$ such that the combinatorial Laplacian matrix in Eqs. (62-63) can be expressed as $\mathbf{L}^{\left(q\right)}_{C}=\mathbf{F}^{\left(q+1\right)}\left[\mathbf{F}^{\left(q+1\right)}\right]^{\top}+\left[\mathbf{F}^{\left(q\right)}\right]^{\top}\mathbf{F}^{\left(q\right)}$. Matrix $\mathbf{S}^{\left(q\right)}$ is a diagonal matrix that discretizes the Hodge star [60] on simplicial complexes. The Hodge star is an operator that transforms $q$-forms to $\left(h-q\right)$-forms in differential geometry theories [60]. Meanwhile, the original Hodge Laplacian, $\Delta^{\left(q\right)}_{D}$, can be approximated by a discrete counterpart, $\left[\mathbf{S}^{\left(q\right)}\right]^{-1}\mathbf{L}^{\left(q\right)}_{C}$ [60].
@@
Our idea arises from a simple fact. No matter we implement a SRG framework with the multi-order Laplacian or the high-order path Laplacian, when this SRG renormalizes the system on a $p$-order according to the properties of $q$-order interactions ($p\leq q$), we can always record the set of units in $\mathbf{G}^{\left(p\right)}_{k}$ that are aggregated into each macro-unit in $\mathbf{G}^{\left(p\right)}_{k+1}$. By repeating this recording in every iteration, we can recursively find the set of initial units in $\mathbf{G}^{\left(p\right)}_{1}$ aggregated into each macro-unit in $\mathbf{G}^{\left(p\right)}_{k+1}$ for each $k\geq 1$. In other words, we can know how the initial structure formed by system units is changed by the SRG after $k$ iterations. If we set $p=1$ and search simplices in the initial system, $\mathbf{G}^{\left(1\right)}_{1}$, we can use the recursively tracked information to identify how many simplices in $\mathbf{G}^{\left(1\right)}_{1}$ are added, removed, or modified by the SRG after $k$ iterations. Therefore, similar to the filtration process in the persistent homology theory [62, 76], the SRG also offers an opportunity to study the lifetime of different topological properties (e.g., the Betti number). Given a simplicial complex $K$, the SRG also creates a filtration sequence of sub-complexes $\left(K_{k}\right)_{k=0}^{m}$ (here $m$ denotes the number of iterations in the SRG). Certainly, the SRG is different from the common filtration processes that persistently add or remove simplices following deterministic rules [62, 76]. It has more intricate effects on the topological properties of the system.
Appendix D The SRG in a form of conventional renormalization group theories
@@
where $\left[\mathbf{L}^{\left(q\right)}\right]_{ij}$ can be selected as $\left[\mathbf{L}^{\left(q\right)}_{M}\right]_{ij}$ or $\left[\mathbf{L}^{\left(q\right)}_{H}\right]_{ij}$. We refer to $\mathcal{H}^{\left(q\right)}$ as the $q$-order functional Hamiltonian, which characterizes the system behaviours on the $q$-order. A more strict form of $\mathcal{H}^{\left(q\right)}$ is shown in Eq. (70), where we indicate that the linear sum over all possible pairs of $\left(i,j\right)$ in the first term of Eq. (69) is not really a linear combination of independent items.
If we write $\mathbf{K}_{ij}^{\left(q\right)}=\left[\mathbf{L}^{\left(q\right)}\right]_{ij}+r\delta_{ij}$, we notice that $\mathbf{K}_{ij}^{\left(q\right)}$ is essentially equivalent to the term of $\phi^{4}$ theory analyzed in Ref. [31]. The main difference lies in that the PCA-like RG proposed by Ref. [31] does not study high-order interactions (the key difference created by introducing high-order interactions is highlighted by Eq. (70) in our subsequent analysis). Given this close relation, we can have following ideas after learning Ref. [31]:
@@
      As shown by the perturbation analysis of the PCA-like RG [31], the behaviours of an interacting system during renormalization are not necessarily governed by the central limit theorem even though the coarse graining procedure only considers Gaussian fluctuations. Instead, more non-trivial behaviours are possible to occur during renormalization. This property makes the PCA-like RG as well as our framework be applicable to more complex systems and identify non-Gaussian behaviours [31]. For instance, the original probability distribution to renormalize in diverse complex systems can be non-Gaussian (although the explicit theoretical expression of this initial probability distribution may be unknown to us). Meanwhile, the departures from a Gaussian distribution can become either more or less important under the scale transformation by the RG. These conditions are all possible to create non-trivial phenomena.
@@
      Most importantly, perhaps a frequently neglected point is that $\mathbf{K}^{\left(q\right)}$ can be defined with $q>1$ (in fact, this is the most common case in the SRG). When $q>1$, the term $\mathbf{x}\left(i\right)\mathbf{K}^{\left(q\right)}_{ij}\mathbf{x}\left(j\right)$ can not be treated as the conventional pairwise relation between units $v_{i}$ and $v_{j}$ that only depends on these two units due to the following two reasons:
      • (b1)
        First, extra units are involved in the relation between units $v_{i}$ and $v_{j}$ when $q>1$. Because the system is analyzed on the $q$-order, this relation between units $v_{i}$ and $v_{j}$ implicitly require the participation of $q+1$ units in total (i.e., there are $q-1$ extra units required by this relation). If there is no enough unit satisfying this requirement, it is impossible for units $v_{i}$ and $v_{j}$ to share a $q$-order interaction since the associated $q$-simplex can not be formed. Therefore, although $\mathbf{x}\left(i\right)\mathbf{K}^{\left(q\right)}_{ij}\mathbf{x}\left(j\right)$ seems to have a dyadic form, it is not equivalent to the classic pairwise relation that only involves with units $i$ and $j$.
      • (b2)
        Second, the off-diagonal elements of $\mathbf{K}^{\left(q\right)}$ can be non-linearly coupled with each other when $q>1$.
        • (b2-1)
          Before further analysis, we recall that a conventional linear combination of multiple items, such as $w=x+y+z$, usually requires the combined items to be independent from each other. For instance, if we reduce $x$ by $k$, we are expected to see that $w$ is reduced by $k$ as a direct consequence of our action and other items (e.g., $y$ and $z$) remain unaffected. In an opposite case where some items are coupled with each other, we can not treat $x+y+z$ as a conventional linear combination. For instance, we assume that reducing $x$ by $k$ will also reduce $y$ by $\frac{x-k}{x+z/k}$ since $x$, $y$, and $z$ are coupled by certain non-linear mechanisms. In this situation, the action of reducing $x$ by $k$ does not directly leads us to $w-k=\left(x-k\right)+y+z$. Instead, it triggers a series of subsequent changes and we need to consider all these changes to ensure that the equality still holds.
        • (b2-2)
          We recall this basic property because Eq. (69) includes a sum over non-linearly coupled items. Let us consider a simple instance with $q=3$. We assume that there exists a $3$-simplex formed by units $a$, $b$, $c$, $d$ in the system (i.e., these units share $3$-order interactions). If we change $\mathbf{K}^{\left(3\right)}_{ab}$ by deleting the interaction between units $v_{a}$ and $v_{b}$ on $1$-order, we also affect the values of $\mathbf{K}^{\left(3\right)}_{ac}$, $\mathbf{K}^{\left(3\right)}_{ad}$, $\mathbf{K}^{\left(3\right)}_{bc}$, $\mathbf{K}^{\left(3\right)}_{bd}$, and $\mathbf{K}^{\left(3\right)}_{cd}$ because the original $3$-simplex inevitably collapses after removing the interaction between units $v_{a}$ and $v_{b}$ (i.e., the definition of clique simplex shown in Appendix A is no longer satisfied). In fact, these changes will propagate over units to modify other $3$-simplices and higher-order simplices until all the simplices involving units $v_{a}$ and $v_{b}$ are changed. These cascade changes after a single modification create non-linear dependencies among the summed items in Eq. (69). Meanwhile, they make the analyzed freedom degrees discontinuous (i.e., since some freedom degrees appear and disappear together). To highlight these dependencies, we introduce a more formal definition of $\mathcal{H}^{\left(q\right)}$
          $\mathcal{H}^{\left(q\right)}$ $\simeq\frac{1}{2}\sum_{\sigma^{\left(q\right)}\in K}\sum_{i,j\in\sigma^{\left(q\right)}}h\left(\sigma^{\left(q\right)}/\{i,j\},q\right)\Bigg{[}\mathbf{x}\left(i\right)\mathbf{K}^{\left(q\right)}_{ij}\mathbf{x}\left(j\right)\Bigg{]}-J\sum_{i}\mathbf{x}\left(i\right),$ (70)
          where we treat the system as a simplicial complex, $K$. The first sum goes through every $q$-simplex, $\sigma^{\left(q\right)}$, in the system and the second sum goes through each pair of units, $v_{i}$ and $v_{j}$, in this $q$-simplex. We denote $\sigma^{\left(q\right)}/\{i,j\}$ as the set of all other units in this $q$-simplex apart from units $v_{i}$ and $v_{j}$. The non-linear map $h\left(\cdot,q\right)$ is referred to as the response function on the $q$-order, which checks through each pair of units in $\sigma^{\left(q\right)}/\{i,j\}$ to identify potential changes. We define $h\left(\sigma^{\left(q\right)}/\{i,j\},q\right)=1$ if and only if there is no change on $\sigma^{\left(q\right)}/\{i,j\}$. If there is any change on any pair of units $v_{k}$ and $v_{l}$ in $\sigma^{\left(q\right)}/\{i,j\}$ (e.g., their interactions are removed on any $p$-order and $p\leq q$), we use $h\left(\sigma^{\left(q\right)}/\{i,j\},q\right)\in\left[0,\infty\right)/\{1\}$ to reflect the cascade effects of this change on $\sigma^{\left(q\right)}$. In a special case where units $v_{i}$ and $v_{j}$ only participate in a single $q$-order interaction (i.e., the one described by $\sigma^{\left(q\right)}$) and at least one pair of units in $\sigma^{\left(q\right)}/\{i,j\}$ become decoupled on the $p$-order ($p\leq q$), we have $h\left(\sigma^{\left(q\right)}/\{i,j\},q\right)=0$ because the only $q$-simplex containing $v_{i}$ and $v_{j}$ has collapsed. Unfortunately, we find it difficult to derive an explicit and universal expression of $h\left(\cdot,q\right)$. This function is highly structure-dependent (i.e., determined by the actual properties of $K$). Defining the response function $h\left(\cdot,q\right)$ on representative systems remains as a meaningful question for future studies, which may suggest a new perspective to study the cascade effects of perturbations using topology.
        • (b2-3)
          In brief, there are numerous non-linear coupling relations between $\mathbf{x}\left(i\right)\mathbf{K}^{\left(q\right)}_{ij}\mathbf{x}\left(j\right)$ for different pairs of $\left(i,j\right)$. In a sense, these coupling relations are the reflections of the un-decomposable properties of high-order interactions. Although we can write $\mathcal{H}^{\left(q\right)}$ in the form of Eq. (69) to relate it with conventional renormalization group theories, we can not neglect these coupling relations that enable cascade effects to happen over a large region and distinguish the SRG from classic frameworks. In Eq. (70), we have shown that these un-decomposable properties are implicitly reflected by $\mathcal{H}^{\left(q\right)}$, where analyzing the relation between units $i$ and $j$ in a $q$-simplex, $\mathbf{x}\left(i\right)\mathbf{K}^{\left(q\right)}_{ij}\mathbf{x}\left(j\right)$, also requires us to check the statuses of other units in $\sigma^{\left(q\right)}/\{i,j\}$. In other words, the $q+1$ units contained in $\sigma^{\left(q\right)}$ can not be studied independently. This property is consistent with conventional definitions of high-order interactions based on simplicial complexes [14].
    • (c)
      Another frequently neglected point is that $\mathcal{H}^{\left(q\right)}$ is not the actual Hamiltonian of the interacting system under study. Instead, $\mathcal{H}^{\left(q\right)}$ is a representational Hamiltonian obtained by placing the interacting system on $q$-order alone. In real cases, diverse orders of interactions can coexist in an interacting system to shape its dynamics, which can not be completely described by $\mathcal{H}^{\left(q\right)}$. The actual Hamiltonian of system dynamics may explicitly consist of different polynomial terms to describe high-order interactions, which can be seen in existing frameworks of high-order interactions in statistical and quantum field theories (e.g., $q$-spin interactions [122, 123, 124]) and synchronization models (e.g., $q$-oscillator couplings [125, 126, 127]). In these frameworks, a $q$-order polynomial term is explicitly proposed to describe $q$-unit interactions (e.g., $3$-unit interactions are characterized by $g\left(i,j,k\right)\mathbf{x}\left(i\right)\mathbf{x}\left(j\right)\mathbf{x}\left(k\right)$, where $g\left(i,j,k\right)$ reflects $3$-unit correlations). In the SRG, we use $\mathcal{H}^{\left(q\right)}$ to guide the renormalization of $\mathcal{H}^{\left(p\right)}$ ($p\leq q$). During the renormalization process, we only deal with the properties of the interacting system on two orders, yet there can exist other orders that affect the renormalization (e.g., as the most simple instance, if $p<w<q$, the renormalization of $p$-order interactions guided by $q$-order interactions is affected by $w$-order interactions as well). Therefore, there exist numerous latent factors to provide possibilities for non-trivial behaviours to occur during renormalization.
Appendix E The behaviour of the specific heat